OpenStudy (swissgirl):

Let \( \mathbb{Z_4} \) = {[0],[1],[2],[3]} Define k:\( \mathbb{Z_4} \) ---> \( \mathbb{Z_4} \) as follows: k([x])=[2x+2] By comparing images verify the following equality k(k(x)) = [2] for all x in \( \mathbb{Z_4} \)

OpenStudy (anonymous):

which inequality?

OpenStudy (swissgirl):

sorry equality

hero (hero):

What course is this? Discrete Math?

OpenStudy (swissgirl):

nope but smth similar

hero (hero):

What is the name of your course?

OpenStudy (swissgirl):

@KingGeorge can u check this out? Like I have an answer but it isnt soo formal

OpenStudy (kinggeorge):

What's your answer you have so far?

OpenStudy (swissgirl):

k(k(x)) = k(2x+2)=4x+6 Since k is a function in \( \mathbb{Z_4} \) then by mod 4, 4x+6-y=4k for some integer y and k 4x+6-2=4k 4x+4=4 4(x+1)=4k Since 4 divides 4x+6 therefore 4x+6=2mod4. Hence 4x+6 is in teh equivalence class [2]

OpenStudy (swissgirl):

idk like its just a guess. I wasnt sure what i was exactly supposed to show

OpenStudy (kinggeorge):

That works, but I don't think it's exactly how they were intending you to show it. I think they meant for you to look at \[k(k([0]))\]\[k(k([1]))\]\[k(k([2]))\]\[k(k([3]))\]individually, and then verify that they all equal \([2]\).

OpenStudy (swissgirl):

ohhhhhhhhhhhhhhhhhh

OpenStudy (swissgirl):

y cant i ever figure out what to doooo?

OpenStudy (kinggeorge):

Probably because the problems aren't worded spectacularly well.

OpenStudy (swissgirl):

Thanks KGGGGGGG

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (swissgirl):

waiittt can u show me how u wld plug it in?

OpenStudy (kinggeorge):

Let's use [1] for an example. \[k(k([1]))=k([2\cdot1+2])=k([4])=k([0])=[2\cdot0+2]=2\]

OpenStudy (swissgirl):

ohhhh i seeeee

OpenStudy (swissgirl):

so lets sat k(k([0]))=k([2*0+2])=k([2])=[2*2+2]=[6]=[2]

OpenStudy (kinggeorge):

Right.

OpenStudy (swissgirl):

yayyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy

OpenStudy (swissgirl):

Thankkkss uuuuuu veryyy much for saving my soul

OpenStudy (kinggeorge):

lol :P You're welcome.